OpenStudy (anonymous):

Dudley is going on a road trip from Washington, DC to St. Louis, MO and back in his Batmobile. Luckily, this is exactly a 500 mile drive. He completed the Washington, DC to St. Louis leg of his trip going 50 mph. His speedometer died on the way back, but miraculously, Dudley is a genius and knows that his average speed for the entire trip was 90mph. How fast was Dudley going (in miles per hour) on the way back, assuming he traveled at a constant speed?

OpenStudy (foolaroundmath):

1. find total time. 2. find time taken on the way back. 3. calculate speed on the way back. can you try something along these lines ?

OpenStudy (anonymous):

d_1 = 50(t_1) 250 = 50t_1 t_1 = 5

OpenStudy (foolaroundmath):

t_1 = 5 is time taken from Washington to St. Louis Can you make use of that 90 mph and find out total time ?

OpenStudy (anonymous):

Um, I'm not sure what to do there. I got stuck around here last time I tried.

OpenStudy (foolaroundmath):

average speed = total distance / total time we know av speed = 90 mph we know total distance = 500 miles So, total time = ? Now use total time and t_1 to calculate time taken in returning

OpenStudy (anonymous):

90 = 500/x x = 500/90 x = 50/9

OpenStudy (anonymous):

50/9 - 5 = 5/9

OpenStudy (foolaroundmath):

yeah thats the time taken for the return journey. know you know time for returning, you know the distance from St Louis to Washington. CAn you calculate speed when returning ?

OpenStudy (anonymous):

s = d/t s = 250/(5/9) s = 450

OpenStudy (anonymous):

Alright, thank you! THat's the given answer. I think I just forgot that average speed = total distance/total time.